OpenStudy (anonymous):

multiple choice, Given the function f(x) = 1/3x, find the value of f(-6) 1)3 2)18 3)2 4)-2

OpenStudy (saifoo.khan):

wht do u think will be the answer?

OpenStudy (saifoo.khan):

-1/18

OpenStudy (anonymous):

thankyou!

OpenStudy (saifoo.khan):

youwelcome!

OpenStudy (anonymous):

its -2

OpenStudy (anonymous):

oh! awesome! thankyou!!!

OpenStudy (anonymous):

-1/18

OpenStudy (anonymous):

am i correct?

OpenStudy (saifoo.khan):

y is it -2? >.<

OpenStudy (mattfeury):

if x = -6, \[\frac{1}{3} * \frac{-6}{1}\] that is your answer

OpenStudy (mattfeury):

the 6 is outside the fraction

OpenStudy (mattfeury):

eh 'x'

OpenStudy (saifoo.khan):

1 ----- 3(-6) 1 --- -18

OpenStudy (anonymous):

i am super confused!

OpenStudy (anonymous):

when u write an equation like that write it as f(x)=(1/3)x

OpenStudy (saifoo.khan):

Please rewrite the Eq, Wagner

OpenStudy (anonymous):

the answer is -2